A normal distribution has a mean of 50 and a standard deviation of 6. What is the probability that a value selected at random from this data is in the interval from 50 to 62? Express your answer as a percent rounded to the nearest tenth.
A) 49.5%
B) 41.5%
C) 47.5%
D) 50.5%

Answer by stanbon(75887)

A normal distribution has a mean of 50 and a standard deviation of 6. What is the probability that a value selected at random from this data is in the interval from 50 to 62? Express your answer as a percent rounded to the nearest tenth.
------
z(50) = (50-50)/6 = 0
z(62) = (62-50)/6 = 12/6 = 2
----
P(50< x <62) = P(0< z <2) = 47.7%
-------------
Cheers,
Stan H.
-------------
A) 49.5%
B) 41.5%
C) 47.5%
D) 50.5%